OpenStudy (anonymous):

A catcher gloves a 92-mi/h pitch and brings it to rest in 0.15m. If the force exerted by the catcher is 803 N, what is the mass of the ball? could someone help me on this? I'm stuck

OpenStudy (anonymous):

i think \[a=-v _{0}^{2}/2s\] so \[(F/a)=m\] F is negative hence exerted contrary to motion

OpenStudy (turingtest):

hosien's method is correct the general kinematic to use here is\[v_f^2=v_i^2+2ad\]the final velocity here is zero, and acceleration is negative, so this becomes\[0=v_i^2-2ad\to a=\frac{v_i^2}{2d}\]which we can sub into our formula for force\[F=ma\to m=\frac Fa=\frac{2dF}{av_i^2}\]now plug in the given numbers and solve

OpenStudy (anonymous):

Thank y'all so much, that helped me out.

OpenStudy (turingtest):

I'm glad :D

OpenStudy (anonymous):

i'm too .helping together is our role .